OpenStudy (anonymous):

Please help someone Solve the system of equations by graphing then classify the system 5x - 5y = -15 5y - 5x = 15 what is the solution of the system? select the correct choice below a. the solution is --- type an ordered pair b. there are infinitely many solutions -- c. there is no solution --- Is the system consistent or inconsistent? Are the equations dependent or independent?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

The system is inconcistent

OpenStudy (anonymous):

when one equation is dependent or independent?

OpenStudy (anonymous):

not understanding the equation part so does this mean it is independent if the system is inconsistent

OpenStudy (amistre64):

my ti83 matrix gives me: 1 -1 | 3 0 0 | 1 but 5x -5y = -15 -5x +5y = 15 <- *-1 5x -5y = -15 5x -5y = -15 ------------- they are the same equation

OpenStudy (anonymous):

meaning there is no solution, inconsistent and independent

OpenStudy (amistre64):

there are infinitely many solutions since every point of both lines matches at every concievable place

OpenStudy (amistre64):

i cant recall the vocab for consistent and dependant tho

OpenStudy (anonymous):

thanks a system of equations is consistent if it has at least one solution and equations are dependent if the system has infinitely many solutions so i guess i just answered my question huh that would be consistent and dependent

OpenStudy (anonymous):

amistre64 could you lead me in the right direction on a graph whereabout the equations would be marked

OpenStudy (anonymous):

amistre64 can you help with this one i lost my way somewhere in trying to solve and graph the compound inequality 7x + 8 > or = to - 48 or 5x + 2 > or = to - 8 the solution of the compound inequality is {x|x>=to ---?